OpenStudy (anonymous):

Express the following equation in slope-intercept form: y - 11 = - 5 (x + 8) 2

OpenStudy (solomonzelman):

slope intercept form looks like y=mx+b

OpenStudy (solomonzelman):

Try to expand the parenthesis on the right hand side, and rearrange the equation that you're given.

OpenStudy (solomonzelman):

I just don't get the problem fully, is it\[y - 11 = -5(x + 8) 2\]\[OR\]\[y - 11 = - 5 (x + 8) \]

OpenStudy (solomonzelman):

@y - 11 = - 5 (x + 8) 2, which of the above is your question? (Or perhaps you already have the answer and understand how to do it?)

OpenStudy (solomonzelman):

(oops tagged a wrong guy.)

OpenStudy (anonymous):

is 2 part of the equation ?

OpenStudy (solomonzelman):

Yes, that's what I am asking.